VIBRATIONAL DEPENDENCE OF HYDROGEN-BROADENED HALFWIDTHS OF WATER VAPOR TRANSITIONS

\underline{R. GAMACHE}, Department of Environmental, Earth, and Atmospheric Sciences and Center for Atmospheric Research, University of Massachusetts, Lowell, MA 01854; R. LYNCH, Wentworth Institute of Technology, 550 Huntington Ave., Boston, MA 02135.

The pressure broadened halfwidth of water vapor transitions in the ground, ν_{2}, ν_{1}, and ν_{3} bands are calculated with hydrogen as the perturbing species using the fully complex Robert and Bonamy theory.^{a} ^{b} The assumed intermolecular potential is a combination of electrostatic, Lennard-Jones 6-12 atom-atom (expanded to 4th order), induction, and dispersion terms. The results are compared with the measurements of Brown and Plymate ^{c} and the vibrational dependence of the halfwidths investigated.

aD. Robert and J. Bonamy, ”Short range force effects in semiclassical molecular line broadening calculations”, J. Phys. (Paris)

40,923 (1979).

bR. Lynch, R. R. Gamache, and S. P. Neshyba, ”Fully Complex Implementation of the Robert-Bonamy Formalism: Halfwidths and Line Shifts of H_{2}O Broadened by N_{2},” submitted to J. Chem. Phys. November, 1995.

cL. R. Brown and C. Plymate, in press, JQSRT (1996).